# Attributing Culture-Conditioned Generations to Pretraining Corpora

### v1

@@ 1 Introduction
Our work presents a generation attribution framework that allows researchers to clearly trace culture-conditioned generations to memorization of patterns in pretraining data. Our finding suggests that language models are unable to reliably and evenly recall knowledge about global cultures in downstream generations, and resort to repeating a small set of high-frequency symbols. Results from our work is helpful for mitigating biases in culture generations when combined with unlearning or data augmentation frameworks. We hope that we inspire more works on attributing model performance on pretraining data.
@@ 3.1 Types of Symbol-Culture Associations in Culture-conditioned Generations
Memorized Association is when the symbol and culture co-occur in meaningful context (i.e. pretraining documents relevant to both the symbol and the culture) with high frequency (distinguishable from other cultures) in pretraining corpora and their association is learned by the model naturally. Memorized associations are important and highly desirable because they are grounded on pretraining documents, demonstrating sufficient model memorization of the symbol-culture association during pretraining.
@@
Weak Association Generalization happens for symbols that are neither memorized association with any culture due to insufficient evidence in the pretraining data to confirm strong memorization for them, nor identified as a diffuse association symbol that is broadly generated for the majority of cultures. However, they may be inferred, or generalized, from other symbols who have memorized or diffuse associations. For example, “kimono” is a type of “robe” specific to Japan, even though “robe” is not memorized association with Japan. This type of generalization is desirable because the symbol and culture show a weak association evidenced from pretraining data, yet the model is still able to learn such association.
@@ 3.3 Identifying Knowledge Memorization from Culture-conditioned Generations
We compute contribution score for every culture $C$ in $C_{\textit{G}}$, and we normalize these scores to form a categorical distribution (See examples in Figure 3). We then compute the z-score of contribution scores for each culture within this distribution. Intuitively, if the distribution is flat, then the symbol is not distinguishably associated with any of the cultures; if the distribution is spiked at a few cultures, then these cultures are distinguishable from the rest for their association with the symbol. Therefore, we set the threshold of z-score to 2.6 ($>99.5\%$ of $C_{\textit{G}}$ [^3] https://www.sjsu.edu/faculty/gerstman/EpiInfo/z-table.htm) to find “outliers” in the distribution and classify the symbols as memorized for cultures whose z-score is above the threshold.
@@ 3.4 Analysis on Non-Memorized Associations from Culture-Conditioned Generations
We identify symbols that are generated for at least half of total cultures, i.e. 55 cultures, as being generated out of diffuse association. We count the occurrence frequency of all symbols of diffuse association and all symbols of memorized association in pretraining data using Infinigram API Liu et al. (2024). In order to verify whether large frequency gap causes the model to prioritize diffuse association to memorized association, for each symbol of diffuse association, we calculate its overshadowing ratio as $r=\frac{1}{N}\sum_{j}\;\frac{count(S_{i})}{count(S_{m_{j}})}$, where $count(S_{i})$ is the count of a diffuse association symbol, $count(S_{m_{j}})$ is the count of the $j$-th unique memorized symbol and $N$ is the number of unique memorized symbols.
@@
It occurs when the model generates symbols memorized from one culture for a different culture. We hypothesize that co-occurrence of both cultures may cause memorized associations to generalize. Therefore, as case study, we perform topic modeling on a subset of symbols and cultures. We extract all documents containing both cultures and the generated symbol, and use LDA Blei et al. (2003) and LLAMA-3.1-8B-Instruct Dubey et al. (2024) to extract common topic words in the documents in which the cultures co-occur (Appendix B).
@@
We prompt OLMo-Instruct-7B to generate definitions of memorized in a continued generation task. Then, we map symbols who are not previously categorized as neither memorized or diffuse association symbols to these definitions using F1 score (Appendix C.1). For each symbol who find a mapping with any weak association symbol, the latter is determined to generalize from the former. Please note that such generalization can be cross-cultural in nature: a weak association symbol generated for one culture can as well be traced to memorized symbols of a completely different culture. Some examples of generalizations that are traced to memorized associations re given in Table 2.
To identify symbols that can be traced symbols with diffuse association, we look for generations with symbols that partially contain or are a combination of independent symbols, such as “black t-shirt” or “rice with meat.”
@@ 4.1 Memorization is Limited for Under-Represented Cultures
The limited memorization for under-represented cultures roots in the inadequate representation in the pretraining corpora. According to  Chang et al. (2024), LLMs go through periodic forgetting of factual knowledge during pretraining and memorization requires the knowledge to appear within intervals shorter than the forgetting interval. Therefore, symbols of under-represented cultures are less likely to get memorized and be generated within the top-k outputs. Instead, symbols not belonging to the culture (evidenced by how MEMOed finds insufficient contributory documents) are generated, leading to diffuse association or cross-culture generalization (see Appendix 4.3 and 4.4).
@@ 4.2 Memorized Association Do Not Limit to Culturally-Emblematic Symbols
To dig deeper into the composition of memorized association, we recruit natives from the each respective culture and ask them whether each symbol “originates from” or “is emblematic to” their own culture.
@@ 4.3 Diffuse Association is frequency-driven
We find a moderate-to-high positive correlation for both clothing (Spearman $\rho=0.551$, Kendall $\tau=0.385$) and food (Spearman $\rho=0.519$, Kendall $\tau=0.385$) on overshadowing ratio $r$ (defined in Section 3.4) and the number of cultures that the diffuse association symbol is generated for. This indicates that high pretraining frequency of diffuse association symbols is magnitudes higher than the frequency of memorized symbols, and this increases the chance of diffuse association overshadowing memorized association during sampling. Figure 5 shows that almost all symbols with diffuse association appear at least 1000 times more frequently than symbols with memorized association, for clothing topic.
@@ 4.4 Cross-Culture Generalization from High to Low Frequency Culture
Frequency Analysis. We first observe a strong positive correlation between 1) the culture’s number of topic-related pretraining documents and 2) the frequency of the culture’s memorized symbol being generated for some other cultures (clothing: Spearman $\rho=0.763$, Kendall $\tau=0.574$; food: Spearman $\rho=0.716$, Kendall $\tau=0.531$). Simultaneously, the culture’s number of topic-related pretraining documents is also negatively correlated with the percentage of the culture’s response containing another culture’s memorized symbols (food only: pearman $\rho=-0.521$, Kendall $\tau=-0.364$).
@@ 4.5 Generalization from Weak Association is not Correlated with Memorization
On average, 3.1% and 5.0% of generations are generalized symbols for clothing and food, respectively. Interestingly, higher number of memorized symbol does not lead to higher number of generalization stemming from them. We only see a weak-to-none correlation (Spearman correlation of 0.17 and -0.03 for clothing and food) between the two types of symbols. Table 10 shows the top and bottom 5 cultures for memorized symbols and generalized symbols for the topic food. Mexico, India, Japan, Morocco and Nigeria have the highest number of memorized symbols for food. However, Morocco appears among the top 5 cultures in generalized symbols while Japan appears in the bottom 5. Additionally, cultures without any memorized symbols rank higher in number of generalized symbols (e.g. Yemenis for clothing and Tribagonian for food). Cultures such as these where the model wasn’t able to memorize anything, prompts the need for generalizations in the next token distribution.
@@ 5 Conclusion
In conclusion, our work introduces MEMOed, a framework for attributing culture-conditioned generations of language models to memorization from pretraining data. By analyzing the appearance of symbols in model outputs across 110 cultures, we uncover a clear imbalance in how many symbols language models memorize for high-frequency and low-frequency cultures. In addition, models tend to prioritize generating high-frequency symbols that are not specific to any culture over memorized symbols, while also struggling to generalize from memorized cultural symbols with lower prevalence in the pretraining data. This highlights significant limitations in current pretraining processes, where models prefers frequently occurring, diffusely associated symbols at the expense of diverse, culture-specific knowledge. Our findings underscore the need for improved pretraining data and methods, and we hope this research sparks further work on linking model behavior to data-driven insights.
@@ B.2 Prompt
In Table 5, we extend our study of pre-training documents (Table 3) pertaining to cultural overmemorization from one culture to another for more cases of cultures which generate these memorized symbols with a lower count of relevant documents than the cultures discussed before. We notice suprisingly similiar themes in the pre-training documents such as the discussion around ”religion” in documents where Hijab, Iran and any culture X co-occur. For Kimono and Japan, we notice a similar common theme surrounding ”fashion”. We hypothesize that such common themes also cause models to overmemorize and generate memorizations from one culture into another and not necessarily when the model is devoid of memorizations in the next token space (which would be the case for cultures which have no memorizations).
@@ F.2 Ablation on OLMo-7B Variants
We obtain a moderate-to-strong correlation for both clothing (spearman $\rho=0.591$, Kendall $\tau=0.507$) and food (spearman $\rho=0.829$, Kendall $\tau=0.659$). Compared to OLMo-7B (on 110 cultures) with clothing (spearman $\rho=0.540$, Kendall $\tau=0.421$) and food (spearman $\rho=0.670$, Kendall $\tau=0.507$), we see that even though OLMo-7B-0424 is tested on smaller number of cultures, for both clothing and food, the correlation of OLMo-7B-0424. Therefore, the conclusion still holds that higher pretraining document counts of cultures increase the number of memorized symbols in culture-conditioned generations.
@@ F.3 Ablation on Z-Score for MEMOed
When z=2, we get still get a moderate-to-strong correlation between 1) the number of memorized symbols for a culture and 2) the count of documents in which the culture appears in the pretraining corpora: for clothing, we obtain a spearman correlation of 0.569 and a Kendall correlation of 0.445; food food, we obtain a spearman correlation of 0.688 and a Kendall correlation of 0.519. This correlation is lower but similar to the original correlations found for z=2.6 (food: Spearman=0.670 and Kendall=0.507; clothing: Spearman=0.540 and Kendall=0.421), showing that our conclusion on the relationship between a culture’s memorized symbols and the culture’s frequency in pretraining data is robust to different z-score threshold.
@@ H.1 Cross-Culture Generalization
favorite_music music, song, songs, album, albums, band, bands, singer, singers,
musician, musicians, genre, genres, concert, concerts
music_instrument music instrument, music instruments, instrument, instruments
exercise_routine exercise, routine, workout, sport, sports
favorite_show_or_movie movie, movies, film, films, TV show, TV shows, TV series, cinema
@@
picture picture, pictures, painting, paintings, portrait, portraits
statue statue, statues, sculpture, sculptures

### v2

@@ 1 Introduction
In summary, our work presents a generation attribution framework that allows researchers to clearly trace culture-conditioned generations to memorization of patterns in pretraining data. Our findings suggest that language models are unable to reliably and evenly recall knowledge about global cultures in downstream generations, and resort to repeating a small set of high-frequency symbols. Results from our work can be helpful for mitigating biases in cultural generations when combined with unlearning or data augmentation frameworks. We hope that we inspire more works on attributing model performance to pretraining data.
@@ 3.1 Types of Symbol-Culture Associations in Culture-conditioned Generations
Memorized Association is when the symbol and culture co-occur in meaningful context with high frequency in pretraining corpora (i.e. the co-occurring pretraining documents are relevant to both the symbol and the culture and their count is distinguishable from other cultures) and their association is learned by the model naturally. Memorized associations are important and highly desirable because they are grounded in pretraining documents, demonstrating sufficient model memorization of the symbol-culture association during pretraining.
@@
Weak Association Generalization happens for symbols that are neither identified as a memorized association with any culture due to insufficient evidence in the pretraining data to confirm strong memorization for them, nor identified as a diffuse association symbol that is broadly generated for the majority of cultures. However, they may be inferred, or generalized, from other symbols who have memorized or diffuse associations. For example, “kimono” is a type of “robe” specific to Japan, even though “robe” is not memorized association with Japan. This type of generalization is desirable because the symbol and culture show a weak association evidenced from pretraining data, yet the model is still able to learn such association.
@@ 3.3 Identifying Knowledge Memorization from Culture-conditioned Generations
We compute contribution score for every culture $C$ in $C_{\textit{G}}$, and we normalize these scores to form a categorical distribution (See examples in Figure 3). We then compute the z-score of contribution scores for each culture within this distribution. Intuitively, if the distribution is flat, then the symbol is not distinguishably associated with any of the cultures. However, if the distribution spikes at a few cultures, then these cultures are distinguishable from the rest for their association with the symbol. Therefore, we set the threshold of z-score to 2.6 ($>99.5\%$ of $C_{\textit{G}}$ [^3] https://www.sjsu.edu/faculty/gerstman/EpiInfo/z-table.htm) to find “outliers” in the distribution and classify the symbols as memorized for cultures whose z-score is above the threshold.
@@ 3.4 Analysis on Non-Memorized Associations from Culture-Conditioned Generations
We identify symbols that are generated for at least half of total cultures, i.e. 55 cultures, as being generated out of diffuse association. We count the occurrence frequency of all symbols of diffuse association and all symbols of memorized association in pretraining data using Infinigram API (Liu et al., 2024). In order to verify whether large frequency gap causes the model to prioritize diffuse association to memorized association, for each symbol of diffuse association, we calculate its overshadowing ratio as $r=\frac{1}{N}\sum_{j}\;\frac{count(S_{i})}{count(S_{m_{j}})}$, where $count(S_{i})$ is the count of a diffuse association symbol, $count(S_{m_{j}})$ is the count of the $j$-th unique memorized symbol and $N$ is the number of unique memorized symbols.
@@
It occurs when the model generates symbols memorized from one culture for a different culture. We hypothesize that co-occurrence of both cultures may cause memorized associations to generalize. Therefore, as case study, we perform topic modeling on a subset of symbols and cultures. We extract all documents containing both cultures and the generated symbol, and use LDA (Blei et al., 2003) and LLAMA-3.1-8B-Instruct (Dubey et al., 2024) to extract common topic words in the documents in which the cultures co-occur (Appendix B).
@@
We prompt OLMo-Instruct-7B to generate definitions of memorized symbols in a continued generation task (Appendix C.1). Then, we map symbols who are previously categorized as neither memorized nor diffuse association symbols to these definitions using F1 score. For each symbol mapped with any weak association symbol, the latter is determined to generalize from the former. Please note that such generalization can be cross-cultural in nature: a weak association symbol generated for one culture can as well be traced to memorized symbols of a completely different culture. Some examples of generalizations that are traced to memorized associations are given in Table 2.
To identify symbols that can be traced to symbols with diffuse association, we look for generations with symbols that partially contain or are a combination of diffuse association symbols , such as “black t-shirt” or “rice with meat.”
@@ 4.1 Memorization is Limited for Under-Represented Cultures
The limited memorization for under-represented cultures roots in the inadequate representation in the pretraining corpora. According to  Chang et al. (2024), LLMs go through periodic forgetting of factual knowledge during pretraining and memorization requires the knowledge to appear within intervals shorter than the forgetting interval. Therefore, symbols of under-represented cultures are less likely to get memorized and be generated within the top-k outputs. Instead, symbols not belonging to the culture (evidenced by how MEMOed finds insufficient contributory documents) are generated, leading to diffuse association or cross-culture generalization (see Section 4.3 and 4.4).
@@ 4.2 Memorized Association Do Not Limit to Culturally-Emblematic Symbols
To dig deeper into the composition of memorized association, we recruit natives from each respective culture and ask them whether each symbol “originates from” or “is emblematic to” their own culture.
@@ 4.3 Diffuse Association is frequency-driven
We find a moderate-to-high positive correlation for both clothing (Spearman $\rho=0.551$, Kendall $\tau=0.385$) and food (Spearman $\rho=0.519$, Kendall $\tau=0.385$) on overshadowing ratio $r$ (defined in Section 3.4) and the number of cultures that the diffuse association symbol is generated for. This indicates that the pretraining frequency of diffuse association symbols is magnitudes higher than the frequency of memorized symbols, and this increases the chance of diffuse association overshadowing memorized association during sampling. Figure 5 shows that almost all symbols with diffuse association appear at least 1000 times more frequently than symbols with memorized association in the pre-training data, for the topic clothing.
@@ 4.4 Cross-Culture Generalization from High to Low Frequency Culture
Frequency Analysis. We first observe a strong positive correlation between 1) the culture’s number of topic-related pretraining documents and 2) the frequency of the culture’s memorized symbol being generated for some other cultures (clothing: Spearman $\rho=0.763$, Kendall $\tau=0.574$; food: Spearman $\rho=0.716$, Kendall $\tau=0.531$). Simultaneously, the culture’s number of topic-related pretraining documents is also negatively correlated with the percentage of the culture’s response containing another culture’s memorized symbols (food only: Spearman $\rho=-0.521$, Kendall $\tau=-0.364$).
@@ 4.5 Generalization from Weak Association is not Correlated with Memorization
On average, 3.1% and 5.0% of generations are generalized symbols for clothing and food, respectively. Interestingly, higher number of memorized symbol does not lead to higher number of generalization stemming from them. We only see a weak-to-none correlation (Spearman correlation of 0.17 and -0.03 for clothing and food) between the two types of symbols. Table 10 shows the top and bottom 5 cultures for memorized symbols and generalized symbols for the topic food. Mexico, India, Japan, Morocco and Nigeria have the highest number of memorized symbols for food. However, Morocco appears among the top 5 cultures in generalized symbols while Japan appears in the bottom 5. Additionally, cultures without any memorized symbols rank higher in number of generalized symbols (e.g. Yemenis for clothing and Tribagonian for food). Cultures such as these where the model wasn’t able to memorize anything, prompts the need for generalizations in the next token sampling process.
@@ 5 Conclusion
In conclusion, our work introduces MEMOed, a framework for attributing culture-conditioned generations of language models to memorization from pretraining data. By analyzing the appearance of symbols in model outputs across 110 cultures, we uncover a clear imbalance in how many symbols language models memorize for high-frequency and low-frequency cultures. In addition, models tend to prioritize generating high-frequency symbols that are not specific to any culture over memorized symbols, while also struggling to generalize from memorized cultural symbols with lower prevalence in the pretraining data. This highlights significant limitations in current pretraining processes, where models prefer frequently occurring, diffusely associated symbols at the expense of diverse, culture-specific knowledge. Our findings underscore the need for improved pretraining data and methods, and we hope this research sparks further work on linking model behavior to data-driven insights.
Acknowledgments
This research is supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #2022-22072200006, the Defense Advanced Research Projects Agency with award HR00112220046, and NSF IIS 2048211. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government. We would like to thank all the collaborators in USC INK research lab for their constructive feedback on the work.
@@ B.2 Prompt
In Table 5, we extend our study of pre-training documents (Table 3) pertaining to cross-cultural generalization from one culture to another for more cases of cultures which generate these memorized symbols with a lower count of relevant documents than the cultures discussed before. We notice suprisingly similiar themes in the pre-training documents such as the discussion around ”religion” in documents where Hijab, Iran and any culture X co-occur. For Kimono and Japan, we notice a similar common theme surrounding ”fashion”. We hypothesize that such common themes also cause models to overmemorize and generate memorizations from one culture into another and not necessarily when the model is devoid of memorizations in the next token space (which would be the case for cultures which have no memorizations).
@@ F.2 Ablation on OLMo-7B Variants
We obtain a moderate-to-strong correlation for both clothing (spearman $\rho=0.591$, Kendall $\tau=0.507$) and food (spearman $\rho=0.829$, Kendall $\tau=0.659$). Compared to OLMo-7B (on 110 cultures) with clothing (spearman $\rho=0.540$, Kendall $\tau=0.421$) and food (spearman $\rho=0.670$, Kendall $\tau=0.507$), we see that even though OLMo-7B-0424 is tested on smaller number of cultures, for both clothing and food, the correlation of OLMo-7B-0424is more strongly positive. Therefore, the conclusion still holds that higher pretraining document counts of cultures increase the number of memorized symbols in culture-conditioned generations.
@@ F.3 Ablation on Z-Score for MEMOed
When $z=2$, we still get a moderate-to-strong correlation between 1) the number of memorized symbols for a culture and 2) the count of documents in which the culture appears in the pretraining corpora: for clothing, we obtain a spearman correlation of 0.569 and a Kendall correlation of 0.445; for food, we obtain a spearman correlation of 0.688 and a Kendall correlation of 0.519. This correlation is lower but similar to the original correlations found for z=2.6 (food: Spearman=0.670 and Kendall=0.507; clothing: Spearman=0.540 and Kendall=0.421), showing that our conclusion on the relationship between a culture’s memorized symbols and the culture’s frequency in pretraining data is robust to a different z-score threshold.